# Atomic sheaves on hyper-Kähler manifolds via Bridgeland moduli spaces

### v2

@@ 3.4. Bayer–Macrì divisors
for any morphism $C\to\mathcal{M}^{Y}_{\sigma_{Y}}(a,b)$ from an integral curve $C$, where $\mathcal{E}_{C}$ is the associated family and $T_{Y_{C}}:=T_{Y}\otimes\mathrm{id}_{\mathrm{D}^{\mathrm{b}}(C)}$ is the base change of $T_{Y}$ [^5] The functor $T_{Y_{C}}:=T_{Y}\otimes\mathrm{id}_{\mathrm{D}^{\mathrm{b}}(C)}\colon\mathrm{D}^{\mathrm{b}}(Y_{C})\to\mathrm{D}^{\mathrm{b}}(Y_{C})$ is defined by the pull-back of the kernel of the composition $\mathrm{D}^{\mathrm{b}}(Y)\xrightarrow{\mathrm{pr}_{Y}}\mathcal{K}u(Y)\hookrightarrow\mathrm{D}^{\mathrm{b}}(Y)$ to $\mathrm{D}^{\mathrm{b}}(Y_{C}\times_{C}Y_{C})$.. Note that $\mathcal{E}_{C}\in\mathrm{D}^{\mathrm{b}}(X_{C})$ by [8, Lemma 8.3(1)]. From [43, Lemma 2.41], we obtain an isomorphism $p_{Y*}(T_{Y_{C}}(\mathcal{E}_{C}))\cong T_{Y}(p_{Y*}(\mathcal{E}_{C}))$. This implies
@@
as $T_{Y}$ acts trivially on $\mathrm{K}_{\mathrm{num}}(\mathcal{K}u(Y))$ by [25, Lemma 4.10].
@@
then (11) follows from the base change of [25, Equation (35)]. This completes the proof of (6). Finally, the ampleness of $\ell_{\sigma_{Y}}$ follows from the facts that $\ell_{\sigma_{X}}$ is ample (cf. [69, Theorem 3.1]) and the morphism $r$ is finite according to Theorem 3.7. ∎
@@
Similarly, using [25, Equation (40)] instead of [25, Equation (35)], we have the following result for GM fivefolds.
@@ 4.1.2. Atomic objects
Such a vector $\tilde{v}(E)\in\tilde{\mathrm{H}}(X,\mathbb{Q})$ is called an extended Mukai vector of $E$ (cf. [6, Definition 4.16] and [5, Proposition 3.3]). In [54, Section 6], the $1$-dimensional subspace of $\tilde{\mathrm{H}}(X,\mathbb{Q})$ spanned by $\tilde{v}(E)$ is called the LLV line of $E$.
@@
By our assumption, $\mathcal{X}_{b}$ is smooth, hence $\mathcal{X}_{b}\hookrightarrow\mathcal{X}$ is a locally complete intersection. Moreover, $\mathrm{codim}_{\mathcal{Y}}(\mathcal{Y}_{b})=\mathrm{codim}_{\mathcal{X}}(\mathcal{X}_{b})=\dim B$. Then using [43, Corollary 2.27], we obtain
@@
As in [5] and [54, Lemma 6.10], for a sheaf $E\in\operatorname{\mathrm{Coh}}(X)$, we have
@@ 4.2. Atomic sheaves supported on Lagrangian subvarieties
By [54, Theorem 6.24] and [5, Proposition 3.11], each torsion atomic sheaf is supported on points or Lagrangian subvarieties. We make the following definition.
@@
is the projection. Since $\chi_{i_{*}\operatorname{\mathcal{O}}_{L}}$ is of rank one, it follows that $\chi_{i_{*}\operatorname{\mathcal{O}}_{M_{t}}}$ has rank at most one for each $t=1,2$. Then the result follows from the non-triviality of $\chi_{i_{*}\operatorname{\mathcal{O}}_{M_{t}}}$ as in the arguments in [54, Lemma 3.9, Remark 3.10] and [5, Section 7.2]. ∎
@@ 4.3. A criterion
To this end, we only need to generalize [5, Proposition 7.2] to our setting. Indeed. the argument there applies analogously without changes. We denote $\mathrm{ker}(\chi^{\mathrm{coh}}_{i_{*}[L]})$ by $K$. Because the image of $i$ is Lagrangian, we see $\mathrm{H}^{2}(X,\operatorname{\mathcal{O}}_{X})\subset K$. Moreover, by the existence of the $\mathfrak{sl}_{2}$-triple $(e_{\omega},h_{\omega},\Lambda_{\omega})\subset\mathfrak{g}(X)_{\mathbb{C}}$ (cf. [72, Section 2]) where $\omega\in\mathrm{H}^{0}(X,\Omega_{X}^{2})$ is the holomorphic symplectic 2-form, the same calculation as in [5, Proposition 7.2] gives $\mathrm{H}^{0}(X,\wedge^{2}\mathcal{T}_{X})\subset K$.
@@
From the proof of Theorem 4.10, we can directly deduce the following criterion:
@@
Let $X$ be a projective hyper-Kähler manifold and $i\colon L\to X$ be an immersed Lagrangian submanifold such that the map $i^{*}\colon\mathrm{H}^{2}(X,\mathbb{Q})\to\mathrm{H}^{2}(L,\mathbb{Q})$ has rank one. Then $i_{*}E$ is atomic for any $E\in\mathrm{D}^{\mathrm{b}}(L)$ satisfying $v(i_{*}E)\exp(\lambda)=i_{*}[L]$ for some $\lambda\in\mathrm{H}^{1,1}_{\mathbb{Q}}(X)$.
@@
The proof of [77, Lemma 1.5] adapts to our situation without changes. It suffices to prove this for cohomology with real coefficients. By the adjunction formula, we have an inclusion $\ker(i^{*})\subset\ker(i_{*}[L]\cup-)$. To show the reverse inclusion, as in the proof of [77, Lemma 1.5], we choose a Kähler class $\lambda\in\mathrm{H}^{2}(X,\mathbb{R})$ such that $i^{*}\lambda$ is also a Kähler class. Now it suffices to prove that the intersection form
@@
is non-degenerate after restricting it to $\mathrm{im}(i^{*})\subset\mathrm{H}^{1,1}_{\mathbb{R}}(L)$. Then the result follows from the Hodge index theorem as in [77, Lemma 1.5] since $i^{*}\lambda$ is a Kähler class. ∎
@@
Let $\omega$ be the holomorphic symplectic $2$-form of $X$. We denote by $\mathcal{N}^{\vee}_{L/X}$ the $-1$ degree cohomology sheaf of the cotangent complex $\mathbb{L}_{L/X}$ of $i$. When $i$ is a closed embedding, $\mathcal{N}^{\vee}_{L/X}$ is exactly the conormal bundle of $L$ in $X$. Since $i$ is a local complete intersection morphism by [71, Tag 0E9K], $\mathcal{N}^{\vee}_{L/X}$ is a bundle (cf. [71, Tag 0FK2] and [71, Tag 068B]). Using the standard exact triangle of cotangent complexes by [71, Tag 0E44], we obtain exact sequences
@@
As $\omega$ is a holomorphic symplectic $2$-form, it induces an isomorphism $\omega\colon\mathcal{T}_{X}\xrightarrow{\cong}\Omega_{X}$. To adapt the argument in [5, Lemma 7.4], it suffices to prove that there exists an isomorphism $\mathcal{N}_{L/X}\cong\Omega_{L}$. According to [33, Remark 4.4], there is a natural map
@@ 5.2. Fixed loci of involutions are $1$-obstructed
First, we assume that $X$ is non-Hodge-special. As in [67, Section 8.5], up to an equivalence of Kuznetsov components, it is enough to prove the statement when $X$ is a non-Hodge-special special GM fourfold. Let $j\colon Y\hookrightarrow X$ be the branch divisor of $X$, which is an ordinary GM threefold. By [25, Proposition 5.3], for any Serre-invariant stability condition $\sigma_{Y}$ on $\mathcal{K}u(Y)$, the functor $\mathrm{pr}_{X}\circ j_{*}$ induces a morphism
@@
Next, we assume that $X$ is an arbitrary smooth GM fourfold. By [66, Theorem 4.18], we may assume that $X$ is general in the sense that the canonical quadric surface in $X$ is smooth (cf. [66, Remark 2.2]). Therefore, using Theorem 3.2, we can find a smooth connected quasi-projective curve $C$ and a family $\mathcal{X}\to C$ of GM fourfolds with a full triangulated subcategory $\mathcal{K}u(\mathcal{X})\subset\mathrm{D}^{\mathrm{b}}(\mathcal{X})$ such that $\mathcal{K}u(\mathcal{X})_{c}\simeq\mathcal{K}u(\mathcal{X}_{c})$ for each $c\in C$. Additionally,
@@
From [13, Lemma 5.9], there is an auto-equivalence $T_{\mathcal{X}}$ on $\mathcal{K}u(\mathcal{X})$ such that $(T_{\mathcal{X}})|_{c}\simeq T_{\mathcal{X}_{c}}$ for each $c\in C$. By [13, Proposition 5.7] and the proof of [67, Theorem 8.2], such equivalence induces an automorphism $\tau_{\mathcal{X}}\colon M^{\mathcal{X}}_{\underline{\sigma}}(a,b)\to M^{\mathcal{X}}_{\underline{\sigma}}(a,b)$, which is an involution since it is an involution at each fiber over $C$ and $M^{\mathcal{X}}_{\underline{\sigma}}(a,b)$ is smooth. Therefore, the fixed locus $\mathrm{Fix}(\tau_{\mathcal{X}})$ is smooth projective over $C$ and satisfies $(\mathrm{Fix}(\tau_{\mathcal{X}}))|_{c}\cong\mathrm{Fix}(\tau_{\mathcal{X}_{c}})$ for each $c$. By Proposition 5.1, the restriction map
@@
If $Y$ is an ordinary GM fivefold, by [42, Corollary 6.5], we can find an ordinary GM threefold $Y^{\prime}$ with an equivalence $\mathcal{K}u(Y)\simeq\mathcal{K}u(Y^{\prime})$. Thus, we only need to prove the statement when $Y$ is an ordinary GM threefold. As $Y$ is ordinary, it is the branch divisor of a special GM fourfold $X$.
Let $\sigma_{X}\in\operatorname{Stab}^{\circ}(\mathcal{K}u(X))$ be a generic stability condition on $\mathcal{K}u(X)$ with $Z_{\sigma_{X}}(\Lambda_{i})=Z_{\sigma_{Y}}(\lambda_{i})$ for each $i=1,2$. By [25, Proposition 5.3], the functor $\mathrm{pr}_{X}\circ j_{*}$ induces a morphism
@@ 5.3. Families of immersed atomic Lagrangians
Let $p\colon M^{\mathcal{Y}}_{\underline{\sigma^{\prime}}}(a,b)\to C$ and $q\colon M^{\mathcal{X}}_{\underline{\sigma}}(a,b)\to C$ be the relative moduli spaces, which are proper. We denote by $r_{c}\colon M^{\mathcal{Y}_{c}}_{\underline{\sigma^{\prime}}_{c}}(a,b)\to M^{\mathcal{X}_{c}}_{\underline{\sigma}_{c}}(a,b)$ the morphism in Theorem 3.7 for each $c\in C$. By assumption, the push-forward of $\mathcal{Y}\to\mathcal{X}$ and the relative projection functor $\mathrm{D}^{\mathrm{b}}(\mathcal{X})\to\mathcal{K}u(\mathcal{X})$ induce a morphism $r^{\prime}\colon M^{\mathcal{Y}}_{\underline{\sigma^{\prime}}}(a,b)\to M^{\mathcal{X}}_{\underline{\sigma}}(a,b)$ such that $r^{\prime}_{c}=r_{c}$. Therefore, $r^{\prime}$ is proper and quasi-finite, hence finite. Moreover, by [25, Proposition 5.1(1)], the fibers of $p$ are equidimensional of the same dimension and $M^{\mathcal{Y}}_{\underline{\sigma^{\prime}}}(a,b)$ is also equidimensional. By Proposition 4.4, this implies that $r_{0*}\operatorname{\mathcal{O}}_{M^{Y}_{\sigma_{Y}}(a,b)}$ is atomic because $r_{1*}\operatorname{\mathcal{O}}_{M_{\underline{\sigma}_{1}}^{Y^{\prime}}(a,b)}$ is atomic. ∎

### v3

@@ 3.4. Bayer–Macrì divisors
for any morphism $C\to\mathcal{M}^{Y}_{\sigma_{Y}}(a,b)$ from an integral curve $C$, where $\mathcal{E}_{C}$ is the associated family and $T_{Y_{C}}:=T_{Y}\otimes\mathrm{id}_{\mathrm{D}^{\mathrm{b}}(C)}$ is the base change of $T_{Y}$ [^7] The functor $T_{Y_{C}}:=T_{Y}\otimes\mathrm{id}_{\mathrm{D}^{\mathrm{b}}(C)}\colon\mathrm{D}^{\mathrm{b}}(Y_{C})\to\mathrm{D}^{\mathrm{b}}(Y_{C})$ is defined by the pull-back of the kernel of the composition $\mathrm{D}^{\mathrm{b}}(Y)\xrightarrow{\mathrm{pr}_{Y}}\mathcal{K}u(Y)\hookrightarrow\mathrm{D}^{\mathrm{b}}(Y)$ to $\mathrm{D}^{\mathrm{b}}(Y_{C}\times_{C}Y_{C})$.. Note that $\mathcal{E}_{C}\in\mathrm{D}^{\mathrm{b}}(X_{C})$ by [8, Lemma 8.3(1)]. From [47, Lemma 2.41], we obtain an isomorphism $p_{Y*}(T_{Y_{C}}(\mathcal{E}_{C}))\cong T_{Y}(p_{Y*}(\mathcal{E}_{C}))$. This implies
@@
as $T_{Y}$ acts trivially on $\mathrm{K}_{\mathrm{num}}(\mathcal{K}u(Y))$ by [27, Lemma 4.10].
@@
then (11) follows from the base change of [27, Equation (35)]. This completes the proof of (6). Finally, the ampleness of $\ell_{\sigma_{Y}}$ follows from the facts that $\ell_{\sigma_{X}}$ is ample (cf. [74, Theorem 3.1]) and the morphism $r$ is finite according to Theorem 3.7. ∎
@@
Similarly, using [27, Equation (40)] instead of [27, Equation (35)], we have the following result for GM fivefolds.
@@ 4.1.2. Atomic objects
Such a vector $\tilde{v}(E)\in\tilde{\mathrm{H}}(X,\mathbb{Q})$ is called an extended Mukai vector of $E$ (cf. [6, Definition 4.16] and [5, Proposition 3.3]). In [59, Section 6], the $1$-dimensional subspace of $\tilde{\mathrm{H}}(X,\mathbb{Q})$ spanned by $\tilde{v}(E)$ is called the LLV line of $E$. Note that by definition, the atomicity of an object $E$ is detected by the linear span of $v(E)$.
@@
By our assumption, $\mathcal{X}_{b}$ is smooth, hence $\mathcal{X}_{b}\hookrightarrow\mathcal{X}$ is a locally complete intersection. Moreover, $\mathrm{codim}_{\mathcal{Y}}(\mathcal{Y}_{b})=\mathrm{codim}_{\mathcal{X}}(\mathcal{X}_{b})=\dim B$. Then using [47, Corollary 2.27], we obtain
@@
As in [5] and [59, Lemma 6.10], for a sheaf $E\in\operatorname{\mathrm{Coh}}(X)$, we have
@@ 4.2. Atomic sheaves supported on Lagrangian subvarieties
By [59, Theorem 6.24] and [5, Proposition 3.11], each torsion atomic sheaf is supported on points or Lagrangian subvarieties. We make the following definition.
@@
is the projection. Since $\chi_{i_{*}\operatorname{\mathcal{O}}_{L}}$ is of rank one, it follows that $\chi_{i_{*}\operatorname{\mathcal{O}}_{M_{t}}}$ has rank at most one for each $t=1,2$. Then the result follows from the non-triviality of $\chi_{i_{*}\operatorname{\mathcal{O}}_{M_{t}}}$ as in the arguments in [59, Lemma 3.9, Remark 3.10] and [5, Section 7.2]. ∎
@@ 4.3. A criterion
To this end, we only need to generalize [5, Proposition 7.2] to our setting. Indeed. the argument there applies analogously without changes. We denote $\mathrm{ker}(\chi^{\mathrm{coh}}_{i_{*}[L]})$ by $K$. Because the image of $i$ is Lagrangian, we see $\mathrm{H}^{2}(X,\operatorname{\mathcal{O}}_{X})\subset K$. Moreover, by the existence of the $\mathfrak{sl}_{2}$-triple $(e_{\omega},h_{\omega},\Lambda_{\omega})\subset\mathfrak{g}(X)_{\mathbb{C}}$ (cf. [77, Section 2]) where $\omega\in\mathrm{H}^{0}(X,\Omega_{X}^{2})$ is the holomorphic symplectic 2-form, the same calculation as in [5, Proposition 7.2] gives $\mathrm{H}^{0}(X,\wedge^{2}\mathcal{T}_{X})\subset K$.
@@
From the proof of Theorem 4.10, we can directly deduce the following criterion.
@@
Let $X$ be a projective hyper-Kähler manifold and $i\colon L\to X$ be an immersed Lagrangian submanifold such that the map $i^{*}\colon\mathrm{H}^{2}(X,\mathbb{Q})\to\mathrm{H}^{2}(L,\mathbb{Q})$ has rank one. Then $i_{*}E$ is atomic for any $E\in\mathrm{D}^{\mathrm{b}}(L)$ satisfying $v(i_{*}E)\exp(\lambda)\in\mathbb{Q}\cdot i_{*}[L]$ for some $\lambda\in\mathrm{H}^{1,1}_{\mathbb{Q}}(X)$.
@@
The proof of [82, Lemma 1.5] adapts to our situation without changes. It suffices to prove this for cohomology with real coefficients. By the adjunction formula, we have an inclusion $\ker(i^{*})\subset\ker(i_{*}[L]\cup-)$. To show the reverse inclusion, as in the proof of [82, Lemma 1.5], we choose a Kähler class $\lambda\in\mathrm{H}^{2}(X,\mathbb{R})$ such that $i^{*}\lambda$ is also a Kähler class. Now it suffices to prove that the intersection form
@@
is non-degenerate after restricting it to $\mathrm{im}(i^{*})\subset\mathrm{H}^{1,1}_{\mathbb{R}}(L)$. Then the result follows from the Hodge index theorem as in [82, Lemma 1.5] since $i^{*}\lambda$ is a Kähler class. ∎
@@
Let $\omega$ be the holomorphic symplectic $2$-form of $X$. We denote by $\mathcal{N}^{\vee}_{L/X}$ the $-1$ degree cohomology sheaf of the cotangent complex $\mathbb{L}_{L/X}$ of $i$. When $i$ is a closed embedding, $\mathcal{N}^{\vee}_{L/X}$ is exactly the conormal bundle of $L$ in $X$. Since $i$ is a local complete intersection morphism by [76, Tag 0E9K], $\mathcal{N}^{\vee}_{L/X}$ is a bundle (cf. [76, Tag 0FK2] and [76, Tag 068B]). Using the standard exact triangle of cotangent complexes by [76, Tag 0E44], we obtain exact sequences
@@
As $\omega$ is a holomorphic symplectic $2$-form, it induces an isomorphism $\omega\colon\mathcal{T}_{X}\xrightarrow{\cong}\Omega_{X}$. To adapt the argument in [5, Lemma 7.4], it suffices to prove that there exists an isomorphism $\mathcal{N}_{L/X}\cong\Omega_{L}$. According to [37, Remark 4.4], there is a natural map
@@ 5.2. Fixed loci of involutions are $1$-obstructed
First, we assume that $X$ is non-Hodge-special. As in [72, Section 8.5], up to an equivalence of Kuznetsov components, it is enough to prove the statement when $X$ is a non-Hodge-special special GM fourfold. Let $j\colon Y\hookrightarrow X$ be the branch divisor of $X$, which is an ordinary GM threefold. By [27, Proposition 5.3], for any Serre-invariant stability condition $\sigma_{Y}$ on $\mathcal{K}u(Y)$, the functor $\mathrm{pr}_{X}\circ j_{*}$ induces a morphism
@@
Next, we assume that $X$ is an arbitrary smooth GM fourfold. By [71, Theorem 4.18], we may assume that $X$ is general in the sense that the canonical quadric surface in $X$ is smooth (cf. [71, Remark 2.2]). Therefore, using Theorem 3.2, we can find a smooth connected quasi-projective curve $C$ and a family $\mathcal{X}\to C$ of GM fourfolds with a full triangulated subcategory $\mathcal{K}u(\mathcal{X})\subset\mathrm{D}^{\mathrm{b}}(\mathcal{X})$ such that $\mathcal{K}u(\mathcal{X})_{c}\simeq\mathcal{K}u(\mathcal{X}_{c})$ for each $c\in C$. Additionally,
@@
From [14, Lemma 5.9], there is an auto-equivalence $T_{\mathcal{X}}$ on $\mathcal{K}u(\mathcal{X})$ such that $(T_{\mathcal{X}})|_{c}\simeq T_{\mathcal{X}_{c}}$ for each $c\in C$. By [14, Proposition 5.7] and the proof of [72, Theorem 8.2], such equivalence induces an automorphism $\tau_{\mathcal{X}}\colon M^{\mathcal{X}}_{\underline{\sigma}}(a,b)\to M^{\mathcal{X}}_{\underline{\sigma}}(a,b)$, which is an involution since it is an involution at each fiber over $C$ and $M^{\mathcal{X}}_{\underline{\sigma}}(a,b)$ is smooth. Therefore, the fixed locus $\mathrm{Fix}(\tau_{\mathcal{X}})$ is smooth projective over $C$ and satisfies $(\mathrm{Fix}(\tau_{\mathcal{X}}))|_{c}\cong\mathrm{Fix}(\tau_{\mathcal{X}_{c}})$ for each $c$. By Proposition 5.1, the restriction map
@@
If $Y$ is an ordinary GM fivefold, by [46, Corollary 6.5], we can find an ordinary GM threefold $Y^{\prime}$ with an equivalence $\mathcal{K}u(Y)\simeq\mathcal{K}u(Y^{\prime})$. Thus, we only need to prove the statement when $Y$ is an ordinary GM threefold. As $Y$ is ordinary, it is the branch divisor of a special GM fourfold $X$.
Let $\sigma_{X}\in\operatorname{Stab}^{\circ}(\mathcal{K}u(X))$ be a generic stability condition on $\mathcal{K}u(X)$ with $Z_{\sigma_{X}}(\Lambda_{i})=Z_{\sigma_{Y}}(\lambda_{i})$ for each $i=1,2$. By [27, Proposition 5.3], the functor $\mathrm{pr}_{X}\circ j_{*}$ induces a morphism
@@ 5.3. Families of immersed atomic Lagrangians
Let $p\colon M^{\mathcal{Y}}_{\underline{\sigma^{\prime}}}(a,b)\to C$ and $q\colon M^{\mathcal{X}}_{\underline{\sigma}}(a,b)\to C$ be the relative moduli spaces, which are proper. We denote by $r_{c}\colon M^{\mathcal{Y}_{c}}_{\underline{\sigma^{\prime}}_{c}}(a,b)\to M^{\mathcal{X}_{c}}_{\underline{\sigma}_{c}}(a,b)$ the morphism in Theorem 3.7 for each $c\in C$. By assumption, the push-forward of $\mathcal{Y}\to\mathcal{X}$ and the relative projection functor $\mathrm{D}^{\mathrm{b}}(\mathcal{X})\to\mathcal{K}u(\mathcal{X})$ induce a morphism $r^{\prime}\colon M^{\mathcal{Y}}_{\underline{\sigma^{\prime}}}(a,b)\to M^{\mathcal{X}}_{\underline{\sigma}}(a,b)$ such that $r^{\prime}_{c}=r_{c}$. Therefore, $r^{\prime}$ is proper and quasi-finite, hence finite. Moreover, by [27, Proposition 5.1(1)], the fibers of $p$ are equidimensional of the same dimension and $M^{\mathcal{Y}}_{\underline{\sigma^{\prime}}}(a,b)$ is also equidimensional. By Proposition 4.4, this implies that $r_{0*}\operatorname{\mathcal{O}}_{M^{Y}_{\sigma_{Y}}(a,b)}$ is atomic because $r_{1*}\operatorname{\mathcal{O}}_{M_{\underline{\sigma}_{1}}^{Y^{\prime}}(a,b)}$ is atomic. ∎